OpenStudy (anonymous):

lim x approaches 0 (1-x)e^x-1/(x^2)

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(1-x)e^{x-1}}{x^2}\]

OpenStudy (freckles):

is that right?

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(1-x)e^x-1}{x^2} \] or \[\lim_{x \rightarrow 0}(1-x)e^x-\frac{1}{x^2}\]

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

the first one?

OpenStudy (freckles):

what happens if you plug in 0? what do you get?

OpenStudy (anonymous):

The one above the word or

OpenStudy (anonymous):

0 but my answer says -1/2

OpenStudy (freckles):

you shouldn't get 0 when you plug in 0

OpenStudy (freckles):

\[\frac{(1-0)e^{0}-1}{0^2}=\frac{e^0-1}{0^2}=\frac{1-1}{0}=\frac{0}{0}\] you get 0/0 which means you have more work to do

OpenStudy (freckles):

if you did have 0/something other than 0 than yeah it would be 0 but 0/0 means you have more work

OpenStudy (freckles):

do you know l'hospital?

OpenStudy (anonymous):

no

OpenStudy (freckles):

do you know this: \[\lim_{x \rightarrow 0} \frac{e^x-1}{x}=1\]

OpenStudy (anonymous):

For real?

OpenStudy (freckles):

i'm just trying to get a feel for where you are at so i can no how to assist l'hosptial is the only way I'm seeing right now doesn't mean there is algebraic way though

OpenStudy (freckles):

or isn't

OpenStudy (anonymous):

cool I just don't understand how -1/2 is my answer

OpenStudy (zarkon):

use the taylor expansion for \(e^x\)

OpenStudy (anonymous):

What's that?

OpenStudy (freckles):

i think i have another way

OpenStudy (freckles):

let u=x^2 \[\lim_{u \rightarrow 0}\frac{[(1-\sqrt{u})e^{\sqrt{u}}-1]-[(1-\sqrt{0})e^{\sqrt{0}}-1]}{u-0}\] but I wonder if i should consider both cases sqrt(x)=u and also -sqrt(x)=u

OpenStudy (freckles):

\[=[(1-\sqrt{u})e^\sqrt{u}-1]'|_{u=0}\]

OpenStudy (freckles):

and probably need to consider both cases because that right there is actually u approaches 0 from right

OpenStudy (freckles):

i was just using the definition of derivative if you are confused :p

OpenStudy (anonymous):

Oh no wonder

OpenStudy (freckles):

\[\lim_{u \rightarrow 0}\frac{f(u)-f(0)}{u-0}=f'(0)\]

OpenStudy (freckles):

you should find both right and left derivative though they will both be the same